OpenStudy (anonymous):

The revenue function in terms of the number of units sold ,x, is given as R=370 x - 0.3 x^2 where R is the total revenue in dollars. Find the number of units sold x that produces a maximum revenue? Please explain?

OpenStudy (anonymous):

1) draw a picture for \[R = -0.3x^2 +370x=-0.3x\left(x-\frac{3700}{3}\right)\] You probably need the vertex equation 

OpenStudy (anonymous):

Explanation: If you can factor the expression so that it is only made up of products y= (x-?)(x-?) then you know that when each parenthesis is equal to zero you have y=0. So when x=0, (-.3x)=0 and therefore r(x)=0 When x=3700/3 then (3700/3 -3700/3) = 0 and therefore y=0 Parabolas are symmetric, so since x=0 and x=3700/3 have the same y, you can argue that the x value between them (3700/3)/2 should be where the vertex is. Then plug x=(3700/3)/2 into R(x) = -0.3x(x-(3700/2)) for R